OpenStudy (anonymous):

The height of a triangle is 5 m less than its base. The area of the triangle is 42 m2. Find the length of the base. A. 11 m B. 12 m C. 8 m D. 7 m

OpenStudy (anonymous):

@pinkbubbles

OpenStudy (aric200):

Do you know the formula to solve the area of a triangle

OpenStudy (anonymous):

hold on one sec

OpenStudy (anonymous):

@pinkbubbles is going to help me you guys should help some other people

OpenStudy (anonymous):

guys its ok :D

OpenStudy (aric200):

Area = (1/2)* base*height

OpenStudy (anonymous):

thank you @aric200

OpenStudy (anonymous):

ok now we can get down to business

OpenStudy (anonymous):

Area = (1/2)* base*height 42 = (1/2)*x*(x-5)

OpenStudy (anonymous):

84 = x^2 - 5x

OpenStudy (anonymous):

now we have to factorize, x^2 - 12x + 7x - 84 = 0 x(x-12) + 7(x-12) = 0 (x-12)*(x+7) = 0

OpenStudy (anonymous):

so its D

OpenStudy (anonymous):

no

OpenStudy (anonymous):

12m

OpenStudy (anonymous):

oh i was going to say 11 but yea your right